OpenStudy (anonymous):

Factorize \(x^2 - 3y^2 - 8z^2 + 2xy + 14yz + 2zx\)

OpenStudy (cwrw238):

rearrange x^2 + 2xy - 3y^2 -8z^2 + 2zx + 14yz = (x - y)(x + 3y) - 2z(4z - 7t + x)

OpenStudy (anonymous):

That doesn't look nice :( How to continue?

OpenStudy (cwrw238):

yes you are right that should be 7y not 7t but other than that this does not look like the way to go

OpenStudy (cwrw238):

this probably will factor into 2 trinomials..

OpenStudy (anonymous):

\[ x^2+2 x y+2 x z-3 y^2+14 y z-8 z^2=(x+3 y-2 z) (x-y+4 z) \]

OpenStudy (anonymous):

I'm actually thinking this, but it will be EXTREMELY UGLY! \[(A_1x+B_1y+C_1z)(A_2x+B_2y+C_2z)\]\[=A_1A_2x^2+B_1B_2y^2+C_1C_2z^2\]\[+(A_1B_2+A_2B_1)xy+(B_1C_2+B_2C_1)yz+(A_1C_2+C_1A_2)xz\] And solve those A1, A2, ...C1, C2..

OpenStudy (anonymous):

@eliassaab Would you mind showing how you did it?

OpenStudy (anonymous):

Apply your method. You know for sure that A1=A2=1 You can also do some guessing about C1 and C2 since C1C2=-8 and so fourth.

OpenStudy (anonymous):

Also B1 B2=-3

OpenStudy (cwrw238):

aah !

OpenStudy (anonymous):

Oh.. I'm stupid! :( Thanks!!

OpenStudy (anonymous):

no .. you're not ..

OpenStudy (anonymous):

So, A1 = 1 A2 = 1 (Since C1C2 = -8 and C1+C2 = 2) C1 = -2 C2 = 4 When B1 =\(\pm\) 1 or B1 = \(\pm\)3 Then B2 = \(\mp\) 3 or B1 = \(\mp\)1 When B1 = 1 and B2 = -3 B1C2 + C1B2 4 + 6 = 10 (rejected) When B1 =-1 and B2 = 3 B1C2 + C1B2 = -4 - 6 = -10 (rejected) When B1 = 3 and B2 = -1 B1C2 + C1B2 12 + 2 = 14 So, B1 = 3 and B2 = -1 Therefore, x^2 - 3y^2 - 8z^2 + 2xy + 14yz + 2zx = (x+3y-2z)(x-y+4z) Wow! Thanks again!!

OpenStudy (anonymous):

yes ..